If you flip a coin, the possible outcomes are HHH, HHT, HTH, HTT, THH, THT, TTH, TTT.
a. What is the probability of getting at most one head?
b. What is the probability of getting at most two tails?

Answer by Fombitz(32388)

a.)At most one head means one head or no head.
4 outcomes:HTT THT TTH TTT
P(at most 1 H)=4/8=1/2

b.)At most two tails means two tails, one tail or no tail. Or simpler, not three tails.
Since only TTT has 3 tails, then all the rest have at most 2 tails.
P(at most 2 T)=7/8